and what if her probability to fall ill is 2/3 while the probability of remaining in good health is 1/3...............let her utility function be given as U(y)=Y powered1/2? then?

(a)will sita accept the insurance?

(b)what is the maximum amountt that sita pay for insurance?
1
Expert's answer
2015-05-21T09:18:08-0400
Future earnings - Rs. 100, p = 1/3­
If she falls ill - Rs. 25, p = 2/3­
U(y) =y1/2­
(a) Sita may accept the insurance, if the probability of illness is too high and the insurance payment is affordable. As p = 2/3, Sita will accept the insurance.
(b) The maximum amount that Sita would pay for the insurance can be calculated according to probabilities:
i(max) = 100*1/3 + 25*2/3 = 33.33 + 16.67 = 50